### v1

# Iterative Trajectory Exploration for Multimodal Agents

Abstract
Multimodal agents, which integrate a controller (e.g., a large language model) with external tools, have demonstrated remarkable capabilities in tackling complex tasks. However, existing agents need to collect a large number of expert data for fine-tuning to adapt to new environments. In this paper, we propose an online self-exploration method for multimodal agents, namely SPORT, via step-wise preference optimization to refine the trajectories of agents, which automatically generates tasks and learns from solving the generated tasks, without any expert annotation. SPORT operates through four iterative components: task synthesis, step sampling, step verification, and preference tuning. First, we synthesize multimodal tasks using language models. Then, we introduce a novel search scheme, where step sampling and step verification are executed alternately to solve each generated task. We employ a verifier to provide AI feedback to construct step-wise preference data. The data is subsequently used to update the controller’s policy through preference tuning, producing a SPORT Agent. By interacting with real environments, the SPORT Agent evolves into a more refined and capable system. Evaluation in the GTA and GAIA benchmarks shows that the SPORT Agent achieves 6.41% and 3.64% improvements, underscoring the generalization and effectiveness introduced by our method. The project page is https://SPORT-Agents.github.io.
Machine Learning, ICML
1 Introduction
Leveraging large language models (LLMs) or vision-language models (VLMs) as controllers to call external tools has become a promising direction in building multimodal agents (Surís et al., 2023; Gupta & Kembhavi, 2023; Zhong et al., 2023; Wang et al., 2024a), achieving impressive performance across multiple complex downstream tasks, such as GUI control (Cheng et al., 2024), multimodal reasoning (Gao et al., 2024a), and embodied AI (Fan et al., 2024). To enhance the planning and reasoning abilities of agents in new environments, existing studies focus on collecting training data to tune the controller of an agent, where commonly used techniques include supervised fine-tuning (SFT) (Hu et al., 2024; Gao et al., 2024b) and reinforcement learning (RL) (Deng et al., 2024c; Xiong et al., 2024b) for alignments with desired behaviors.
The key to agent tuning is collecting sufficient expert data (e.g., tasks, labels, and trajectories to solve tasks). SFT methods collect expert data via human annotation or distillation from closed-source API to fine-tune the controllers. RL methods use the expert data to tune reward models, update policy models (Qi et al., 2024; He et al., 2024), or derive performance data for optimization (Zhang et al., 2024; Putta et al., 2024). However, collecting high-quality expert data may be difficult in new environments. It is labor-extensive and high-cost, and such pre-collected data may lead to biased distributions inconsistent with the target environments, causing inferior results.
Figure 1: Pipeline of the proposed SPORT method, including four iterative components: task generation, step sampling, step verification, and preference tuning.
In this paper, we focus on the multimodal reasoning tasks, where the agents are required to call diverse tools (e.g., web search, visual reasoning, file understanding, and object localization) to answer given questions. We explore whether multimodal agents can improve their performance via online self-exploration without any expert data. We are inspired by existing research in LLMs and VLMs, which has shown impressive performance in self-instruction (Wang et al., 2023; Liu et al., 2023), self-verification (Madaan et al., 2024; Yu et al., 2024), and self-learning (Deng et al., 2024b; Kumar et al., 2024). Based on the above observation, we expect that the agent automatically generates tasks, searches for possible trajectories to call tools in solving synthetic tasks, evaluates these trajectories by itself, and updates controllers using this data. In this case, agents will improve their generalization capability by interacting with environments.
To achieve this goal, we must address three key challenges in such an online self-exploration framework for multimodal reasoning tasks. (1) Lack of off-the-shelf tasks and expert trajectories. There is no off-the-shelf dataset with ground truth annotation for multimodal reasoning tasks. It is non-trivial to identify whether a task is correctly solved, thus making it challenging to select correct trajectories. (2) Difficulty in reward modeling. Unlike other settings (e.g., GUI setting), where sufficient data or pre-defined rules make reward modeling feasible, predicted answers or trajectories in multimodal reasoning cannot be easily and reliably verified. (3) Low sampling efficiency and high cost. Sampling trajectories often involves expensive tools (e.g., LLM APIs, web search), which results in both high monetary and computational costs, making it challenging to scale up.
To solve the above challenges, we propose SPORT, an iterative self-exploration method via step-wise preference optimization to refine trajectories of multimodal agents, as shown in Figure 1. SPORT operates through four iterative components: task synthesis, step sampling, step verification, and preference tuning. First, we generate queries and multimodal files for task synthesis based on provided task seeds. Second, we introduce a new search scheme that samples step-level candidate solutions to call tools. Third, we employ a multimodal verifier that, given the task context, intermediate reasoning states, and step candidates, provides AI-generated feedback to estimate step-level preferences. Finally, we perform step-wise preference tuning to refine the controller’s policy and obtain the SPORT Agent, which is then used to guide trajectory sampling in the next iteration.
SPORT enables agents to autonomously generate tasks and explore tool-use trajectories, removing reliance on expert data or pre-collected datasets. For pre-trained LLMs, providing step-level preference signals is easier than accurate rewards, circumventing the difficulties of reward modeling in complex multimodal tasks. Furthermore, SPORT improves the utilization of sampled trajectories—even for some failed trajectories, we can still discover useful step-level preference data, allowing us to acquire more effective data utilization with the same number of sampled trajectories. These capabilities collectively support stable, scalable self-exploration for complex multimodal tasks.
We conduct experiments on the two multimodal reasoning benchmarks: GTA and GAIA, and results show that our SPORT Agent outperforms the SFT Agent by 6.41% and 3.64%, respectively. This indicates that our SPORT Agent method leads to a more powerful reasoning and planning capability for tool usage by interacting with the environment.
In summary, our contributions are three-fold.
(1) We propose an online self-exploration framework for multimodal agents, through which the agents can adapt to new environments without any expert annotation.
(2) We propose SPORT, an online self-exploration method that utilizes step-wise optimization and AI feedback, providing a possible way to self-learning in complex multimodal environments.
(3) The obtained SPORT Agent achieves significant performance improvements compared with SFT based agents on two popular benchmarks: GTA and GAIA.
2 Related Work
2.1 Agent Tuning
Due to the disparity between the LLMs and the requirements of agents, agent tuning is necessary to adapt to practical tasks in new environments. Research for agent tuning could be divided into two categories: supervised fine-tuning (SFT) and reinforcement learning (RL). SFT methods collect expert trajectory data via distillation from closed-source API (e.g., GPT-4o) (Gao et al., 2024b; Liu et al., 2024d; Zeng et al., 2023) or human annotation (Liu et al., 2024c; Deng et al., 2024a). Then they use these collected data to tune the controller via SFT. However, the SFT methods suffer from huge costs and inferior generalization (Shi et al., 2024). To solve this issue, researchers have paid attention to RL agent tuning methods that allow agents to interact with the environment and learn from the feedback. Some methods utilize the policy gradient technique (Zhou et al., 2024a) to update the controller with a reward model that is designed as a fine-tuned model (Qi et al., 2024; Zhai et al., 2024b), environment feedback (Bai et al., 2024; Zhai et al., 2024a), human-designed rules (He et al., 2024; Zhou et al., 2024b), or tree search results (Deng et al., 2024c). To simplify this procedure, the direct preference optimization (DPO) method is applied to agent tuning (Xiong et al., 2024b), where trajectory-level (Zhang et al., 2024; Song et al., 2024) or step-level preference data (Putta et al., 2024; Chen et al., 2024b) is constructed to update the controller. The above DPO methods construct step-wise preference data based on whole correct trajectories. Nevertheless, such expert or correct trajectories are difficult to obtain in complex multimodal tasks, and the predefined correct trajectories may lead to biased distributions. In contrast, our online self-exploration framework does not rely on any correct trajectories. In addition, it uses an online manner to avoid biased distributions.
2.2 Step-wise Preference Tuning
Preference tuning methods rely on paired data, which is not readily available for complex tasks with multi-step reasoning, making it non-trivial to determine which trajectory is better. Furthermore, for long trajectories, only an overall preference verification can not capture the relationships among steps and ignores the fine-grained preference between different steps. To overcome this problem, step-wise preference has been studied. STEP-DPO (Lai et al., 2024) and SCDPO (Lu et al., 2024) collect step-wise preference data by localizing error steps or disturbing the correct path. OREO (Chen et al., 2024a) and SVPO (Wang et al., 2024b) train value models for step-wize verification and inference guidance. SDPO (Kong et al., 2025) combines step-level, turn-level, and session-level preference data for full-grained optimization. The above methods mainly focus on code generation and math reasoning tasks that are easy to obtain correct trajectories to construct step-wise preference data. In contrast, this method focuses on multimodal agents, where obtaining correct trajectories is challenging. Thus, our agent explores the environment by itself via an online manner, which uses designed AI feedback to construct preference data without any annotation.
2.3 Learning from AI Feedback
Using models to generate AI feedback (natural languages, scores, or preference) for performance improvement has emerged as a critical paradigm (Bai et al., 2022). Existing methods can be broadly divided into three categories based on how to use the obtained AI feedback. Methods of the first category add the AI feedback into prompts for in-context learning (Madaan et al., 2024). LLaVA-Cirtic (Xiong et al., 2024a) collects sufficient data to train a model to provide rich multimodal AI feedback. VLM-F (Liao et al., 2024), VolCaNo (Lee et al., 2024a), and Clarify (Lee et al., 2024b) add AI feedback to VLM prompts to address visual hallucinations. CLOVA (Gao et al., 2024a) and CompAgent (Wang et al., 2024f) are representative agents that use AI feedback to refine prompts for better task-solving. Methods of the second category use AI feedback to filter data for supervised fine-tuning. M-STAR (Liu et al., 2024b) shows its effectiveness in visual mathematical reasoning. APIGen (Liu et al., 2024d), MAT (Gao et al., 2024b), and visualagentbench (Liu et al., 2024c) design verifiers to collect trajectories for agents. Methods of the third category use AI feedback for reinforcement learning. They use LLMs or VLMs to produce rewards for policy gradient optimization (Lee et al., ; Fu et al., ; Rocamonde et al., 2023), and several of them produce preference data for the DPO algorithm (Yu et al., 2024). Different from them, our AI feedback is well-designed for evaluating each step in complex scenarios of multimodal agents.
Figure 2: Demonstrations of the search scheme used in the SPORT method. Given a task, the agent samples potential solutions for each step and verifies their qualities in an online manner. Then, we construct the step-wise preference data based on such self-exploration.
3 Method
3.1 Formulation
We opt for the framework of the ReAct agent (Yao et al., 2023) that performs step-by-step reasoning for tool usage for task solving. In each step, based on the input $x_{i}$, the agent outputs a solution $s_{i}$ for tool calling.
$s_{i}^{\star}=\arg\max\pi_{\theta}(s_{i}|x_{i},T),$ (1)
where $\pi_{\theta}$ is the controller (an VLM in our method) of agents with $\theta$ being the parameters, $x_{i}$ is composed of the task (including a query $Q$ in natural language and multimodal files $F$) and the history $h_{i}$ of previous steps, i.e., $x_{i}=\{Q,F,h_{i}\}$. The solution $s_{i}$ consists of the thought $t_{i}$ and code $c_{i}$, $s_{i}^{=}\{t_{i},c_{i}\}$. $T$ denotes available tools, and we follow the work (Gao et al., 2024b) using the same toolkit. In this case, we further rewrite Eq. (1) as
$t_{i}^{\star},c_{i}^{\star}=\arg\max\pi_{\theta}(t_{i},c_{i}|Q,F,h_{i},T),$ (2)
where $t_{i}^{\star}$ and $c_{i}^{\star}$ are thought and code for the $i$-th step, and the history $h_{i}=\{t_{1},c_{1},o_{1},\cdots,t_{i-1},c_{i-1},o_{i-1}\}$ is composed of thought, code, and observation of previous steps.
Agent tuning aims to update $\theta$ to increase the planning and reasoning capabilities of agents in new environments. This paper proposes an online self-exploration method, SPORT, to update $\theta$ via step-wise preference optimization in refining trajectories, as shown in Figure 1. Concretely, SPORT has iterative components: task synthesis, step sampling, step verification, and preference tuning. In one iteration, SPORT first generates some multimodal tasks. For each generated task, SPORT performs step sampling and step verification alternately to construct step-wise preference data. Finally, SPORT uses the step-wise preference data to tune the controller.
3.2 Task Synthesis
Since a multimodal task is composed of a language query and multimodal files, the task synthesis component is divided into query generation and multimodal file generation. We first generate queries and then generate files, rather than the reverse order, since the multimodal files are diverse (such as DOCX, PPTX, XLSX, and PDF are commonly encountered), and it is challenging to construct a diverse file dataset in advance. In addition, tasks are usually based on multiple files instead of only one. First obtaining files and then generating queries may cause weak relevance of files and produce non-natural queries.
To produce diverse and practical queries, we collect seed from the existing method MAT (Gao et al., 2024b) and employ an LLM (GPT-4o-mini in practice) to generate queries. We feed randomly sampled seed queries, used tools, and a designed prompt to the LLM that generates multiple queries once. Adding tool descriptions to the prompt makes GPT-4o mini better understand what queries can be solved, improving their qualities. For each generated query, we prompt GPT-4o mini to output the needed files. If images are needed, we search for source images from off-the-shelf datasets based on similarities. For other files, we prompt GPT-4o mini to generate Python code to produce files.
3.3 Data construction
The two components: step sampling and step verification are performed to construct high-quality preference data, playing key roles in our method. To avoid potential bias issues in constructed data, we introduce an online search scheme, where the step sampling and step verification are executed alternately in each generated task, as shown in Figure 2.
The step-wise preference data is formulated as a triplet $(x_{i},s_{i}^{pre},s_{i}^{dis})$, where $x_{i}=\{Q,F,h_{i}\}$ denotes the input including the query $Q$, files $F$, and the history $h_{i}$ of previous steps, $s_{i}^{pre}$ is the preferred solution in the current step including, and $s_{i}^{dis}$ is the dispreferred solution.
Given a task with the query $Q$ and files $F$, the agent expands the search space for the first step by sampling $n$ solutions $\{s_{1}^{1},s_{1}^{2},\cdots,s_{1}^{n}\}$, including the thought and code $\{t_{1}^{1},c_{1}^{1},\cdots,t_{1}^{n},c_{1}^{n}\}$ from the controller, and execute them to obtain $n$ observations $\{o_{1}^{1},\cdots,o_{1}^{n}\}$. Then, we feed the query $Q$, $n$ solutions, and $n$ observations to an LLM, and ask the LLM to select the best solution $\{t_{1}^{*},c_{1}^{*}\}$ with its corresponding observation $o_{1}^{*}$.
Along $\{t_{1}^{*},c_{1}^{*}\}$ and $o_{1}^{*}$, we expand the search space for the second step. Regarding $\{t_{1}^{*},c_{1}^{*},o_{1}^{*}\}$ as the history $h_{2}$, the controller samples $n$ solutions $\{t_{2}^{1},c_{2}^{1},\cdots,t_{2}^{n},c_{2}^{n}\}$ from the controller, and executes them to obtain $n$ observations $\{o_{2}^{1},\cdots,o_{2}^{n}\}$. We feed the query $Q$, history $h_{2}$, $n$ solutions in this step, and $n$ corresponding observations to an LLM, and ask the LLM to select the best solution $\{t_{2}^{*},c_{2}^{*}\}$ with its corresponding observation $o_{2}^{*}$. In this case, the agent gradually expands the search space and selects the best solution for each step, until the agent believes that the task is over.
Assume there are $m$ steps in solving one task. In this case, we could collect $m(n-1)$ preference data pairs. In the $i$-th step, the selected best solution $\{t_{i}^{*},c_{i}^{*}\}$ is the preferred output, and the rest $n-1$ solutions are the dispreferred outputs, denoted as $\mathcal{D}_{i}^{dis}$, $|\mathcal{D}_{i}^{dis}|=n-1$. The preference data in one task is denoted as $\{(x,\{t_{i}^{*},c_{i}^{*}\},\{t_{i}^{j},c_{i}^{j}\})\}$, where $\{t_{i}^{j},c_{i}^{j}\}\in\mathcal{D}_{i}^{dis}$ and $i\in[1,m]$.
3.4 Preference Tuning
Objective. In one iteration, we may generate multiple tasks and construct preference data for them. After that, we denote the obtained preference data set as $\mathcal{D}=\{(x_{i},s_{i}^{pre},s_{i}^{dis})\}$, and the number of data is $|\mathcal{D}|=d$. We choose the flexible DPO algorithm,
[…]
where $\pi_{\theta}$ is the controller to be updated, $\pi_{ref}$ is the controller for reference (the model after SFT in practice), $\beta$ is the weighting parameter that controls the deviation from the reference controller, and $\sigma(\cdot)$ is the logistic function.
Efficiency. One issue is the potential low efficiency in such an online self-exploration framework. A key hyperparameter is $d$, that is, how much preference data is collected in one iteration. If we set a large $d$, it means that we collect a large number of preference data for agent tuning in one iteration. This may cause the problem of data bias, as the data required by the controller may change during training, while our data is generated using a fixed controller in the beginning. If we set a small $d$, it means that we collect a small number of preference data for agent tuning in one iteration. This may cause low efficiency, since we need to load and release the model weights (controller, parameter gradients, tools, and verifier) frequently. To balance it, we set $d$ as 500 in practice via empirical exploration.
Training Scheme. The online self-exploration is performed after an SFT stage for controllers, since the effectinveness of the self-exploration framework requires the controller to have the ability to generate accurate solutions. The SFT stage is the same as MAT (Gao et al., 2024b), where 20K trajectories are used to align the agent controller (Qwen2VL-7B in practice) with desirable outputs. In the self-exploration stage, we use preference tuning to update Qwen2-VL using 16K step-wise preference data in all, where $d$ is 500 and there are 4 iterations in total. The preference tuning process is summarized in Algorithm 1.
Algorithm 1 Training process in SPORT.
0:  Seeds of multimodal tasks, initial agent controller $\pi_{\theta}$, and $\pi_{ref}=\pi_{\theta}$. Preference data $\mathcal{D}=\emptyset$.
0:  Updated agent controller $\pi_{\theta^{*}}$
1:  while Not converged do
2:     Set $\mathcal{D}=\emptyset$.
3:     Randomly sample task seeds, and send them to an LLM to generate tasks.
4:     for Each generated task do
5:        for the $i$-step in solving the task do
6:           Sample $n$ solutions $\{t_{i}^{1},c_{i}^{1},\cdots,t_{i}^{n},c_{i}^{n}\}$ based on the history $h_{i}$, and execute them to obtain results $\{o_{i}^{1},\cdots,o_{i}^{n}\}$.
7:           Select the best solution $\{t_{i}^{\star},c_{i}^{\star}$}.
8:           Construct $n-1$ preference pairs, and add them into $\mathcal{D}$.
9:           Add $t_{i}^{\star},c_{i}^{\star},o_{i}^{\star}$ into $h_{i}$.
10:        end for
11:     end for
12:     Use $\mathcal{D}$ to update $\pi_{\theta}$ via the preference tuning algorithm in Eq. (3).
13:  end while
4 Experiments
4.1 Setting
The performance of the proposed SPORT approach is assessed on the GTA (Wang et al., 2024c) and GAIA (Mialon et al., 2023) benchmarks. Results are compared against agents powered by both closed-source models (e.g., GPT-4, GPT-4o, Claude3) and open-source models, including LLaMA-3-70B-instruct (Dubey et al., 2024), Qwen1.5-72B-chat (Bai et al., 2023), LLaVA-NeXT-8B (Liu et al., 2024a), InternVL2-8B (Chen et al., 2024c), Qwen2-VL-7B (Wang et al., 2024d), and MiniCPM-V-8.5B (Yao et al., 2024). Specifically, we perform direct comparisons with leading agents, such as Lego Agent (AgentLego Contributors, 2023), Sibyl Agent (Wang et al., 2024e), and Warm-up Act Agent (Mialon et al., 2023). As a baseline, we use the Huggingface Agent (HF Agent) (HuggingFace Contributors, 2024), which operates with the same toolset as the Sport Agent . We first evaluate these agents on two benchmarks, then assess the quality of produced preference data, and finally show several visualization examples to demonstrate the effectiveness of our method.
We employ the Qwen-2-VL model as the controller. In the training process of our VLM controller, we freeze the vision encoder and visual token compressor, and fine-tune the language model using LoRA (Hu et al., 2022). We set the rank as $32$ and apply LoRA on query, key, and value projection matrices in all self-attention layers. We use the AdamW optimizer with a cosine annealing scheduler. The learning rate is $61.0e-$ 6 and the batch size is $2$ per device. We set the max context window as $10240$ to support complex trajectories of our agent.
Benchmark. The GTA and GAIA benchmarks serve as robust evaluation frameworks for assessing multimodal agents. The GTA benchmark includes 229 tasks paired with 252 images, where task completion requires 2 to 8 steps, with most tasks involving 2 to 4 steps. This benchmark challenges multimodal agents to exhibit advanced perception, operational skills, logical reasoning, and creative thinking based on visual data. In real-world multimodal scenarios, agents often need to handle diverse file formats such as PPTX, PDF, and XLSX. To evaluate agent performance on such files, the GAIA benchmark is used, comprising 446 tasks across 109 files. GAIA’s tasks are organized into three levels, with task complexity varying from 2 steps to sequences of indefinite length. It evaluates document comprehension, web navigation, logical reasoning, and summarization abilities.
Metric. Following existing methods (Wang et al., 2024c; Gao et al., 2024b), we assess agent performance using three key metrics: AnsAcc, ToolAcc, and CodeExec for the GTA benchmark. AnsAcc gauges the accuracy of predicted answers. ToolAcc evaluates the correctness of tool usage and the quality of answer summaries. CodeExec measures the percentage of generated code that executes without errors. In the GAIA benchmark, we focus on measuring AnsAcc at its three levels.
4.2 GTA Results
The results on the GTA benchmark are shown in Table 1, where key metrics including AnsAcc, ToolAcc, and CodeExec are reported. Our agent surpasses the Lego agent that utilizes closed-source models (e.g., GPT-4 and GPT-4o), as well as the HF agent that uses closed-source models and open-source models (e.g., InternVL2-8B), showcasing the ability of our SPORT Agent to tackle complex tasks with greater efficiency. A comparison between agents through SFT (i.e., T3-Agent) and our SPORT Agent demonstrates the effectiveness of our online self-exploration framework and the advantages of our Step-wise optimization approach. Our method has about $7\%$ improvements on the final accuracy, since it calls more suitable tools ($8\%$ improvements) and reduces code error ($7\%$ improvements). Compared with the HF agent using GPT-4o and GPT-4o mini, our agent achieves higher ToolAcc and comparable CodeExec. This indicates that the proposed SPORT method improves the planning and reasoning capabilities of agents again.
Table 1: Results on the GTA benchmark
Method Controller AnsAcc ToolAcc CodeExec
Lego Agent GPT-4 46.59 - -
Lego Agent GPT-4o 41.52 - -
Lego Agent GPT-3.5-turbo 23.62 - -
Lego Agent Claude3-opus 23.44 - -
Lego Agent Qwen1.5-72B-chat 13.32 - -
Lego Agent LLaMA3-70B-instruct 8.32 - -
HF Agent GPT-4o 57.05 63.41 95.12
HF Agent GPT-4o mini 57.69 56.10 100.00
HF Agent LLaVA-NeXT-8B 14.10 14.97 25.08
HF Agent InternVL2-8B 32.05 36.75 52.18
HF Agent MiniCPM-V-8.5B 33.97 36.59 56.10
HF Agent Qwen2-VL-7B 42.31 44.85 65.19
T3-Agent MAT-MiniCPM-V-8.5B 52.56 65.85 80.49
T3-Agent MAT-Qwen2-VL-7B 53.85 64.63 84.32
Sport Agent Tuned Qwen2-VL-7B 60.26 72.41 91.87
4.3 Ablation
4.3.1 Effectiveness of iteration step size $d$.
We conduct an ablation study on the GTA benchmark to investigate the impact of the iteration step size $d$. We set $d\in\{200,500,1000\}$, adjusting the number of iterations to (5, 2, 1), respectively, to ensure a total of 1000 tasks are processed in each setting. As shown in Table 2, setting $d=500$ yields the best overall performance across all metrics. When $d=1000$, the model sees all tasks in a single pass, which limits adaptability and leads to a slight drop in answer accuracy and execution success. In contrast, using a smaller step size ($d=200$) results in less diverse updates per iteration, which may reduce robustness. These results suggest that a moderate value of $d$ offers a good balance between update frequency and data diversity, leading to more stable and effective training.
Table 2: Ablation on iteration step $d$ in the GTA benchmark
$\boldsymbol{d}$ AnsAcc (%) ToolAcc (%) CodeExec (%)
200 56.41 68.58 88.46
500 57.69 69.87 89.74
1000 57.05 69.87 88.46
4.3.2 Applying DPO to MAT Preference Data
To quantify the impact of directly optimizing MAT trajectories with preference-based learning, we perform an ablation by applying DPO to the original MAT-SFT model. Specifically, we treat the MM-Traj data (GPT-4o generated) as “preferred” samples and synthetically generate an equal number of “dispreferred” examples via the MAT-SFT model, matching the total volume of our SPORT preference data. We then fine-tune MAT-SFT using DPO under this constructed preference dataset.
Table 3 reports results on the GTA benchmark. Compared to vanilla MAT-SFT, MAT-SFT-DPO yields only modest improvements (AnsAcc +1.28, ToolAcc +2.67, CodeExec +1.58), indicating that naïvely applying DPO to MAT provides limited gains. In contrast, SPORT substantially outperforms MAT-SFT-DPO (AnsAcc +5.13, ToolAcc +5.11, CodeExec +5.97), demonstrating the effectiveness of our framework in leveraging diverse, multimodal preference data.
Table 3: Ablation on preference data: comparison between MAT-SFT, MAT-SFT-DPO, and SPORT on the GTA benchmark.
Method AnsAcc (%) ToolAcc (%) CodeExec (%)
MAT-SFT 53.85 64.63 84.32
MAT-SFT-DPO 55.13 67.30 85.90
SPORT (Ours) 60.26 72.41 91.87
4.4 GAIA Results
In Table 4, we report the performance of Sport Agent on the GAIA validation set. Sport Agent achieves best results among agents that use open-source models, surpassing the best-performing open-source model, Qwen2-VL-7B, about $11\%$ on AnsAcc. The consistent improvements across different levels underscore the efficacy of our online self-exploration framework. Furthermore, Sport Agent demonstrates significant gains over the SFT-tuned controller, with an improvement about $4\%$ over MAT-Qwen2-VL-7B. However, when compared to agents leveraging closed-source models such as GPT-4, Sport Agent exhibits a slight performance gap. We attribute this discrepancy to the larger model sizes and more extensive training data available to closed-source models.
Table 4: Results on the validation set of the GAIA benchmark
Method Controller AnsAcc Level 1 Level 2 Level 3
Sibyl Agent GPT-4-turbo 29.70 43.40 27.90 7.70
Warm-up Act GPT-4-turbo 17.60 30.20 15.10 0.00
HF Agent GPT-4o 33.40 47.17 31.40 11.54
HF Agent GPT-4o mini 26.06 33.96 27.91 3.84
HF Agent LLaVA-NeXT-8B 3.64 9.43 1.16 0.00
HF Agent InternVL2-8B 4.85 7.55 4.65 0.00
HF Agent MiniCPM-V-8.5B 7.27 13.21 5.81 0.00
HF Agent Qwen2-VL-7B 9.70 16.98 8.14 0.00
T3-Agent MAT-MiniCPM-V-8.5B 15.15 26.42 11.63 3.84
T3-Agent MAT-Qwen2-VL-7B 16.97 26.42 15.12 3.84
Sport Agent Tuned Qwen2-VL-7B 20.61 35.85 16.28 3.84
Figure 3: Tool distribution for the chosen and rejected steps.
4.5 Statistic
We aggregated step-wise preference data from all the $4$ iterations, each with $d=500$ and a sampling size of 5, resulting in a total of 16K samples. We analyze the differences between the $chosen$ and $rejected$ step-wise data in terms of code error rate, tool selection, and content variations.
Tool Distribution. We analyze the distribution of tools in the chosen and rejected steps by examining their frequency of occurrence, as shown in Figure 3. The results highlight distinct differences in tool usage between the two groups.
In the chosen steps, tools such as visualizer (2101 occurrences) and objectloc (1051 occurrences) are used more frequently. In contrast, in the rejected steps, objectloc (1442 occurrences) and visualizer (1524 occurrences) are more prevalent. Additionally, tools like ocr and seg appear more often in the rejected steps. To quantify the overall discrepancy in tool usage, we computed a tool distribution difference of $45.62\%$, indicating a substantial variation in the tool selection between the chosen and rejected steps.
Figure 4: Example of constructed preference data.
[…]
Table 5: Comparison of code error rates between chosen and rejected steps.
chosen rejected
Code error rate 18.35% 81.94%
Table 6: Comparison of BLEU scores (lower scores indicate greater discrimination) between our verifier and random selection.
Verifier BLEU-1 ($\downarrow$) BLEU-2 ($\downarrow$) BLEU-3 ($\downarrow$) BLEU-4 ($\downarrow$)
Random Select 0.53 0.41 0.36 0.34
Ours 0.30 0.18 0.14 0.11
Content Difference. We compared the BLEU scores of steps selected by our verifier and those selected randomly. BLEU scores measure the similarity between different sequences, with lower scores indicating higher discrimination. As shown in Table 6, our verifier consistently achieved lower BLEU scores across all n-grams compared to random selection. Specifically, for BLEU-1, BLEU-2, BLEU-3, and BLEU-4, our verifier’s scores were $0.30$, $0.18$, $0.14$, and $0.11$, respectively, while random selection yielded higher scores of $0.53$, $0.41$, $0.36$, and $0.34$. The results demonstrate that our verifier selects more distinct and differentiated steps, enhancing the quality of the chosen steps.
Figure 5: Comparisons between SFT Agent and our SPORT Agent.
Table 7: Average scores from humans.
Task Trajectory Preference
Reasonableness Naturalness Tool Content Code
8.16 8.48 8.78 9.08 8.44 82%
4.6 Data Quality
To evaluate the effectiveness of the constructed preference data, we conducted a user study involving 20 AI researchers with coding and development experience from various universities and research institutes. These participants were not provided with any background information about our methodology; instead, they were only briefed on the purpose and functionality of the agent. They were required to justify whether the preferred and dispreferred pairs were proper.
The evaluation was performed for tasks and trajectories, based on five criteria: (1) ‘Reasonableness’ to evaluate whether the generated tasks are infeasible; (2) ‘Naturalness’ to evaluate whether the generated tasks are natural; (3) ‘Code’ to evaluate the accuracy of code in solutions; (4) ‘Tool’ to assess the appropriateness of tool selection; and (5) ‘Parameter’ to assess the correctness of parameter passing. Participants provided scores ranging from 1 to 10, with higher scores indicating better performance. As shown in Table 7, the average scores for our task and framework exceeded 8, demonstrating the validity of the preference data collected by our approach.
To further assess the effectiveness of the verifier in generating preference data, we conducted an additional study with a separate group of 20 researchers. Each participant was asked to evaluate 50 steps sampled in parallel by making preference selections. The preferences were determined using the same three criteria: code accuracy, tool selection appropriateness, and parameter passing correctness. We measured the agreements between the verifier’s preferences and those of the human participants, denoted as ‘Preference’ in Table 7. The results revealed an $82\%$ overlap. This high level of agreement validates the reliability of the verifier in capturing human-like preferences.
4.7 Visualization
We visualize the preference data generated by the online self-exploration framework, as shown in Figure 4. The verifier can successfully choose solutions that lead to correct intermediate results. In step 1, the solution that produces the correct brand and basic information about the smartphone is selected as the preferred data. In step 2, the solution that searches for contents most relevant to the task is selected as the preferred data. It compares the smartphone with mid-range competitors, while the rest compares the smartphone with a special one.
We visualized the task-solving procedure of the SPORT Agent compared with the SFT Agent (T3 agent), as shown in Figure 5. The SPORT Agent after the step-wise preference tuning can well solve the issues of code hallucination and tool error. For example, in case (a), the SPORT Agent correctly sends the image path to the tool. In case (b), the SPORT Agent utilizes tools to solve the task, instead of producing an answer via hallucination.
5 Conclusion
In this paper, we have presented an online self-exploration framework for multimodal agents, through which the agents can learn via automatic interaction with new environments without accessing any expert data. Based on this framework, we have presented a step-wise optimization for refining trajectories (SPORT), which can produce in-distribution preference data in complex environments. Given proper prompts, the proposed SPORT method can generate diverse multimodal tasks and make good verification of agent solutions aligned with humans. Experiments on two challenging benchmarks: GTA and GAIA, show that the proposed SPORT method achieves significant improvements on multimodal agents, demonstrating the effectiveness.
Limitation. The verifier plays an important role in the current SPORT method. However, it heavily relies on human-designed rules and prompts, causing inferior generalization for some outlines. In the future, we will explore the self-exploration techniques for the verifier, that is, learning to verify, through which the verifier can adapt to new environments with the controller together. Furthermore, we will explore the theoretical guarantee for the verifier, allowing it to scale to open settings.
Impact Statement
This paper presents work whose goal is to advance the field of AI agents. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Comparison with Existing Sampling Frameworks
Our method constructs step-level preference data for complex multimodal tasks. To contextualize its contributions, we compare against representative reinforcement-learning–based sampling frameworks along three dimensions: (1) Task Domain, (2) Collection Granularity, and (3) Annotation Format. Table 8 summarizes this comparison.
Table 8: Comparison with existing sampling schemes
Method Task Domain Collection Granularity Annotation Format
WebRL (Qi et al., 2024) GUI control Trajectory-Level Finetune a reward model
PAE (Zhou et al., 2024a) GUI control Trajectory-Level Use a pre-trained model
ETO (Song et al., 2024) GUI control & Embodied AI Trajectory-Level Expert labels for comparisons
DMPO (Shi et al., 2024) GUI control & Embodied AI Trajectory-Level Finetune a reward model
DigiRL (Bai et al., 2024) GUI control Step-Level Finetune a reward model
TP-LLAMA (Chen et al., 2024b) API calling Step-Level Use expert data
IPR (Xiong et al., 2024b) GUI control & Embodied AI Step-Level Use expert data
StepAgent-Inverse (Deng et al., 2024c) GUI control & Embodied AI Step-Level Use expert data
Ours Multimodal Reasoning Step-Level Use a pre-trained model
The multimodal reasoning domain poses unique challenges that prior sampling frameworks do not adequately address:
  1. 1.
    Data scarcity. There is a shortage of collected tasks and expert trajectories in complex multimodal settings.
  2. 2.
    Inadequate reward modeling. Approaches effective in GUI or API domains—such as pre-trained classifiers or rule-based rewards—struggle to capture multimodal task complexity.
  3. 3.
    Low sampling efficiency. Generating high-quality trajectories for multimodal tasks requires expensive tool usage (e.g. LLM calls, web searches), making naive sampling impractical.
To overcome these issues, we leverage pre-trained LLMs to generate step-wise preference data automatically, providing a scalable and practical solution for training agents on complex multimodal reasoning tasks.
Appendix B Prompt of the Step Verifier
To evaluate the quality of intermediate steps taken by an agent during task execution, we design a step verifier consisting of two key components: a system prompt and a user prompt. The system prompt (Figure 6) provides the verifier model with detailed instructions for evaluating multiple candidate steps (‘CURRENT_STEP‘) based on their coherence, logical progression, and effectiveness in advancing the task. It guides the model to consider contextual alignment with the prior step, tool usage, hallucination, and content relevance. The verifier is required to select the best step and justify its decision in a structured JSON format.
You are an evaluation assistant responsible for analyzing and evaluating agent trajectories. Your goal is to rank < N> ‘CURRENT_STEP‘ entries based on their coherence, logical progression, and effectiveness in addressing the TASK, as observed in the ‘CURRENT_RESULT‘, and their alignment with the ‘PREVIOUS_STEP‘. Input Description: You will receive <N> sets of the following: - ‘PREVIOUS_RESULT‘: The prior results obtained by the agent. - ‘CURRENT_STEP‘: The agent’s output, containing a ‘thought‘ and ‘code‘ intended to complete the task based on the observation. - ‘CURRENT_RESULT‘: The result or state produced by executing the ‘CURRENT_STEP‘. Your Task: 1. Evaluate each ‘CURRENT_STEP‘: - Assess how well the proposed ‘CURRENT_STEP‘ aligns with the context established by the ‘PREVIOUS_STEP‘ and the observation reflected in the ‘CURRENT_RESULT‘. - Check for coherence, logical progression, and contextual relevance. - Prioritize outputs that effectively build upon or adapt to the ‘PREVIOUS_STEP‘ while addressing the ‘CURRENT_RESULT‘. 2. Select the BEST of the ‘CURRENT_STEP‘ entries: - Pick the best ‘CURRENT_STEP‘ according to the following guidelines. 3. Provide a concise explanation for your choice: - Highlight key factors that influenced your decision, such as logical flow, contextual relevance, effectiveness, and uniqueness of the result. Evaluation Guidelines: - Hallucination: Penalize the directly hallucinated content in the code instead of being produced from tools. - Tool selection: Pay attention to whether the controller selects the proper tool. - Best content pass into the tool: For the two step that uses the same tool, pay attention to the query that the controller sends to the tools, such as the ’question’ in visualizer() and ask_search_agent(). - Task Relevance: Ensure the CURRENT_STEP contributes meaningfully to solving the task. - Maintain objectivity and avoid assumptions beyond the provided inputs. Output Format: Return your evaluation in the following JSON structure: { "reason": "<concise_explanation_of_ranking>" "best_id": <An int that indicates the id for the best step. Since there are five CURRENT_RESULTs, the id should only be one of 1,2,3,4, and 5>, }
Figure 6: System prompt for the step verifier.
The user prompt (Figure 7) supplies the concrete input to the verifier, including the task description and a list of candidate step sets. Each step set contains the result from the previous step, the current step (thought and code), and the result produced by executing that step. Together, these prompts simulate a human-like evaluation process, encouraging the model to perform judgment aligned with human preferences in multi-step reasoning scenarios.
The following are the given task, results of previous steps, and result of the current step. TASK: <task> Step Sets: <step_set> # Format of step_set: # { #       ‘PREVIOUS_RESULT‘: <The prior results obtained by the agent.> #       ‘CURRENT_STEP‘: <The agent’s output, containing a ‘thought‘ and ‘code‘ intended to complete the task based on the observation.> #       ‘CURRENT_RESULT‘: <The result or state produced by executing the ‘CURRENT_STEP‘.> # } Now, you need to determine the best of the current steps based on the above information.
Figure 7: User prompt for the step verifier.
Appendix C User Study Interface
C.1 Preference Alignment Study
Figure 8 (a) presents the web interface used to evaluate how well our automated verifier’s preferences align with those of human judges. In each trial, participants were shown a single task case along with a collection of candidate next-step actions (each consisting of a brief “Thought” description and an optional code snippet or tool invocation). These options were the same ones ranked by our verifier, but presented in random order to prevent positional bias.
Participants were instructed to review each candidate step and select the one they considered most appropriate for progressing the task. No additional scoring rubric was provided: judges were simply asked to choose the option they “would use” if they were guiding the model. Once a selection was made, participants clicked “Submit” to lock in their preference and proceed to the next case.
By comparing the human-selected option against the top choice of the verifier, we compute an agreement rate for each model and task type. High agreement indicates that the verifier captures human judgments of step quality; lower agreement reveals areas where the verifier’s ranking diverges from human intuition. These results were aggregated over 50 cases per participant, enabling both per-case analysis and overall statistics on human–verifier alignment.
C.2 Data quality
Figure 8 (b) illustrates the web interface employed in our user study. For each case, participants proceeded through two consecutive scoring phases:
  1. 1.
    Task Evaluation.
    • Reasonableness (1–10): Does the prompt and the displayed interaction trajectory form a logical, feasible, and well-defined user request?
      • 1–3: Highly unreasonable or ill-posed.
      • 4–6: Somewhat reasonable but with noticeable flaws.
      • 7–9: Mostly reasonable, with only minor issues.
      • 10: Fully logical and indistinguishable from genuine user queries.
    • Naturalness (1–10): Is the phrasing realistic, user-like, and fluent?
      • 1–3: Artificial or awkward.
      • 4–6: Acceptable but with unnatural turns of phrase.
      • 7–9: Mostly natural, with only minor awkwardness.
      • 10: Fully natural and conversational.
  2. 2.
    Trajectory Evaluation. After reading both the “Chosen” and “Rejected” responses, participants rated the model’s proposals on three dimensions, each on a 1–10 scale (1 = lowest, 10 = highest):
    • Code Accuracy: Correctness and completeness of any code snippets.
    • Tool Effectiveness: Appropriateness and utility of the suggested APIs or functions.
    • Content Accuracy: Relevance and factual correctness of the text or image descriptions.
In both phases, brief written examples anchored scores at the low, mid, and high ends to promote consistency. Upon completing all six ratings for a case, participants clicked “Next” to submit their responses and proceed to the following item.
subfigureInterface for the user study on verifier performance. Users select their perfered steps.
subfigureInterface for the user study on task execution and trajectory. Users evaluate the quality of tasks and trajectories.
Figure 8: User interface.